### v1

# Sky-Drive: A Distributed Multi-Agent Simulation Platform for Socially-Aware and Human-AI Collaborative Future Transportation

Abstract
Recent advances in autonomous system simulation platforms have significantly enhanced the safe and scalable testing of driving policies. However, existing simulators do not yet fully meet the needs of future transportation research—particularly in modeling socially-aware driving agents and enabling effective human-AI collaboration. This paper introduces Sky-Drive, a novel distributed multi-agent simulation platform that addresses these limitations through four key innovations: (a) a distributed architecture for synchronized simulation across multiple terminals; (b) a multi-modal human-in-the-loop framework integrating diverse sensors to collect rich behavioral data; (c) a human-AI collaboration mechanism supporting continuous and adaptive knowledge exchange; and (d) a digital twin (DT) framework for constructing high-fidelity virtual replicas of real-world transportation environments. Sky-Drive supports diverse applications such as autonomous vehicle (AV)–vulnerable road user (VRU) interaction modeling, human-in-the-loop training, socially-aware reinforcement learning, personalized driving policy, and customized scenario generation. Future extensions will incorporate foundation models for context-aware decision support and hardware-in-the-loop (HIL) testing for real-world validation. By bridging scenario generation, data collection, algorithm training, and hardware integration, Sky-Drive has the potential to become a foundational platform for the next generation of socially-aware and human-centered autonomous transportation research. The demo video and code are available at: https://sky-lab-uw.github.io/Sky-Drive-website/.
Index Terms:
Driving Simulator, Autonomous Vehicles, Human-AI Collaboration, Multi-Agent Simulation, Digital Twin.
I Introduction
Autonomous systems and related technologies have made significant strides in recent years, demonstrating increasing maturity in perception, decision-making, and control capabilities [1, 2, 3, 4]. As these technologies continue to advance, future transportation systems are expected to consist of diverse intelligent agents, including autonomous vehicles (AVs), human-driven vehicles (HVs), delivery robots, flying vehicles, and smart infrastructure [5]. In this emerging ecosystem, each agent must not only ensure its own safe and efficient operation, but also align its behavior with human preferences and societal norms through continuous interactions with various road users, such as pedestrians and cyclists. As a result, future research must move beyond the isolated validation of AV performance toward the broader investigation of social awareness and human-AI collaboration in mixed traffic environments.
Validating autonomous driving technologies in real-world settings presents considerable challenges due to safety risks, limited controllability, and the scale of testing required to demonstrate reliability [6, 7, 8]. To mitigate these barriers, the autonomous driving community has developed a variety of simulation platforms, including CARLA [9], AirSim [10], SUMO [11], Vissim [12], Highway-Env [13], MetaDrive [14], SMARTS [15], CarSim [16] and IPG CarMaker [17]. These platforms have significantly accelerated development by providing controlled testing environments. However, they face key limitations in addressing the evolving needs of future transportation research.
First, while existing simulation platforms can emulate multiple agents on a single machine using rule-based or pre-trained learning-based methods [11, 12, 13], they generally do not support real-time participation of human users (e.g., drivers or pedestrians) across multiple terminals. This limitation hinders the collection of authentic human behavior and human-agent interaction data. Such data is particularly valuable for studying rare but safety-critical scenarios—for example, interactions between AVs, HVs, and pedestrians—which pose significant safety and ethical risks when collected in the real world. A distributed simulation platform that enables participants to assume diverse roles across multiple terminals is urgently needed to safely collect such interaction data and to evaluate interaction algorithms in controlled environments.
Second, existing simulation platforms offer limited support for human-AI collaboration. While they can collect human inputs, these are often treated as low-level control signals rather than high-level feedback for improving autonomous driving algorithms [18, 19, 20]. In contrast, effective human-AI collaboration refers to a bidirectional process in which humans provide feedback not only as commands, but also as indications of preferences, situational understanding, and normative behaviors; AI systems, in turn, assist human drivers by offering real-time guidance, performance feedback, and personalized training. This bidirectional exchange enables AI systems to continuously adapt to human needs and expectations while simultaneously enhancing human driving performance through intelligent support. Additionally, the emergence of foundation models—such as large language models (LLMs) [21, 22] and vision-language models (VLMs) [23]—trained on large-scale, multimodal datasets and equipped with broad world knowledge—offers new opportunities for capturing and utilizing human knowledge [24]. Yet, in most simulators [25, 26, 27], such models are used primarily for scenario generation rather than as active components in human-AI collaborative learning.
Figure 1: Overview of Sky-Drive’s key components and functionalities. (a) a distributed multi-agent architecture enabling synchronized simulation across multiple terminals; (b) a multi-modal human-in-the-loop framework capturing comprehensive behavioral data through integrated sensor systems; (c) a digital twin framework that creates high-fidelity virtual replicas of transportation systems through multi-source data integration; (d) a human-AI collaboration mechanism facilitating knowledge exchange between humans and AI systems; (e) the planned integration of foundation models to enhance decision-making, enabling more adaptive and context-aware human-AI collaboration; (f) a hardware-in-the-loop framework, planned for future integration, ensuring that algorithms are evaluated in real-world environments.
Third, although some simulation platforms have integrated reinforcement learning (RL) capabilities to improve autonomous driving policies [14, 15, 28], they remain largely focused on optimizing vehicle-level metrics such as safety, efficiency, and route completion. However, advancing real-world deployment requires moving beyond individual vehicle performance to incorporate social awareness into the decision-making process. Social awareness refers to an autonomous system’s ability to consider the impact of its actions on surrounding road users and the broader traffic environment [29]. This includes promoting traffic flow stability, enhancing the comfort of other participants, and enabling harmonious interactions between AVs and humans in mixed traffic settings. In this context, transportation science offers a valuable foundation. Decades of research have produced validated traffic flow theories and human behavioral models—such as the intelligent driver model (IDM) [30] and the minimizing overall braking induced by lane changes (MOBIL) [31] model—that can inform the design of socially aware autonomous systems.
To address these challenges, this paper proposes Sky-Drive, an open-source simulation platform designed to advance research in socially aware autonomous driving and human-AI collaboration. Sky-Drive integrates scenario generation, data collection, algorithm training, and hardware integration into a unified platform, supporting distributed multi-agent operation and multi-modal human-in-the-loop interaction. As illustrated in Fig. 1, Sky-Drive introduces four key innovations:
  • Sky-Drive introduces a distributed multi-agent architecture that enables synchronized simulation across multiple devices through a remote procedure call (RPC) networking model. This design allows independent control of agents on separate terminals while maintaining shared environmental states, better reflecting future mixed traffic.
  • Sky-Drive provides a multi-modal human-in-the-loop framework that integrates diverse sensors, including steering wheels, virtual reality (VR) systems, cameras, and smartwatches, to capture rich human behavioral data. A synchronized data processing pipeline correlates these multi-modal streams, enabling detailed analysis of human driving patterns and responses to complex scenarios.
  • Sky-Drive implements an innovative human-AI collaboration mechanism comprising a human as AI mentor (HAIM) module that incorporates human feedback and domain knowledge to guide AI learning, and an AI as human mentor (AIHM) module that provides real-time guidance and personalized training to human drivers.
  • To bridge the gap between simulation and reality, Sky-Drive includes a digital twin (DT) framework that builds high-fidelity virtual replicas of transportation systems by integrating data collected from lab-developed AVs, roadside sensors, traffic cameras, and historical records.
To further enhance Sky-Drive’s capabilities, two major functionalities are planned:
  • Sky-Drive will integrate foundation models at both the system and agent levels. At the system level, foundation models will provide global observation and feedback to optimize simulation dynamics. At the agent level, they will enhance situational understanding and enable safer, more socially aware, and personalized decision-making.
  • Sky-Drive will incorporate a hardware-in-the-loop (HIL) framework via robot operating system (ROS) integration, enabling direct validation of autonomous driving algorithms on physical vehicles and safe evaluation of algorithms without exposing users to real-world risks.
The remainder of this paper is organized as follows: Section II reviews related work in driving simulators. Section III introduces Sky-Drive’s workflow. Section IV details Sky-Drive’s features and technical implementation. Section V demonstrates application examples. Section VI discusses planned future enhancements. Finally, Section VII concludes the paper and outlines future research directions.
II Related Work
II-A Driving Simulators
Driving simulation platforms have evolved significantly to address the growing needs of AVs research. According to Li et al. [32], these simulators can be categorized based on their primary functions and capabilities.
Comprehensive simulators provide end-to-end virtual environments with complete road networks, diverse traffic agents, pedestrians, and detailed sensor models. CARLA [9] and LGSVL [33] represent prominent open-source examples in this category, offering rich environments for testing autonomous driving systems. Commercial solutions such as Nvidia Drive Sim [34] and rFpro [35], alongside academic developments including DeepDrive [36] and GarchingSim [37], provide similar comprehensive capabilities. Another important category is traffic flow simulators, which focus on modeling network-level vehicle movements, traffic congestion, and large-scale traffic scenarios. Notable examples include SUMO [11], Vissim [12], Flow [38], and CityFlow [39]. Recent developments combine SUMO’s traffic modeling with 3D simulators such as CARLA to merge scalability with realism.
Sensory data simulators, such as AirSim [10] and Sim4CV [40], are designed to generate high-fidelity sensor outputs for perception systems. These functionalities are increasingly being integrated into comprehensive simulators while maintaining their critical role in AV perception testing. Driving policy simulators provide configurable environments for evaluating decision-making algorithms. Examples include Highway-Env [13], TORCS [41], SUMMIT [42], MACAD [43], SMARTS [15], and MetaDrive [14]. Additionally, recent data-driven simulators such as Waymax [18], ScenarioNet [19], and Nocturne [20] leverage real-world datasets to generate socially relevant traffic scenarios. Vehicle dynamics simulators, including CarSim [16], IPG CarMaker [17], and Gazebo [44], specialize in accurately modeling vehicle physics, such as suspension responses and tire-road interactions, which are essential for validating control algorithms under realistic conditions.
While existing platforms offer valuable simulation capabilities, certain challenges remain in supporting future transportation research. As shown in Tab. I, most simulators run only on single devices, limiting their ability to model distributed multi-agent scenarios. Additionally, current platforms provide insufficient support for socially-aware algorithms that need to understand complex interactions with diverse road users. Considering CARLA’s established strengths in sensor simulation and visualization, Sky-Drive leverages CARLA as its core engine while extending it with a distributed architecture for synchronized multi-terminal simulation, immersive VR interfaces, and a DT framework. This integration creates an open-source platform specifically designed to support future transportation research.
TABLE I: Comparison of Representative Simulators with Sky-Drive
Distributed Digital Twin Hardware- Traffic Flow AI Framework Human-in-the-
Multi-agent Simulation Environment in-the-Loop Modeling Integration loop Interface
Closed Source
Nvidia Drive Sim [34] - -
rFpro [35] - - -
CarSim [16] - - - -
Matlab [45] - -
Open Source
DeepDrive 2.0 [36] - - - - -
GarchingSim [37] - -
CARLA [9] - -
SUMO [11] - - -
Flow [38] - - - -
CityFlow [39] - - - -
TORCS [41] - - - - -
SUMMIT [42] - - - -
MACAD [43] - - - -
MetaDrive [14] - - -
SMARTS [15] - - - - -
Nocturne [20] - - -
Waymax [18] - - -
Gazebo [44] - - - -
Sky-Drive (Ours)
Note: The “Distributed Multi-agent Simulation” functionality in this table refers to the capability of simulators to synchronize and run multiple agents (e.g., AVs, HVs, and pedestrians) across different computers in real-time simulations. This is distinct from simply running multiple agents concurrently on a single computer, which most simulators can accomplish.
II-B Human–AI Collaboration Environments
Several simulation platforms have contributed to advancing human-AI collaboration in autonomous driving. For instance, NVIDIA’s DRIVE Sim and Omniverse platform [46] support collaboration by generating physics-based synthetic data for training autonomous systems. However, their approach largely enables one-way knowledge transfer—where simulated data informs AI models—without supporting real-time, bidirectional human-AI interaction. Applied Intuition incorporates human-in-the-loop testing to allow operators to validate autonomous decisions, yet its framework is primarily tailored for offline validation rather than continuous learning [47]. MORAI provides digital twin environments that visualize AI decision-making for human drivers, but its interaction remains limited to basic feedback collection without mechanisms for mutual adaptation or learning [48].
More specialized platforms have made progress toward collaborative learning. MIT’s VISTA enables domain adaptation between real and virtual environments, but focuses mainly on perception rather than interactive decision-making [49]. The GAMMA framework introduces mixed-reality traffic incorporating human behavior, though it lacks explicit mechanisms for integrating human expertise into AI learning [50]. Wayve’s LINGO architecture enhances transparency by providing natural language explanations for AI decisions [51], while SafeMod leverages LLMs for bidirectional planning, mimicking human reasoning in autonomous decision-making [52]. Similarly, SurrealDriver generates realistic driving behaviors that align with human expectations using LLMs [53], and DarwinAI’s GenSynth facilitates collaboration between human designers and AI in developing neural networks for driving tasks [54].
Despite these advances, most platforms still fall short in enabling true human-AI knowledge exchange. They often lack mechanisms for the continuous integration of human feedback, resulting in open-loop rather than closed-loop learning processes. Moreover, few platforms support comprehensive multimodal sensing from humans—such as gaze, voice, physiological signals, and control inputs—which are critical for modeling and understanding driving behaviors. Sky-Drive addresses these limitations through its HAIM and AIHM modules, its multi-modal human-in-the-loop framework, and its closed-loop learning architecture that continuously integrate human expertise into AI development.
Figure 2: Workflow of Sky-Drive. (a) scenario generation & data collection through CARLA-based synthetic environments and digital twin integration of real-world traffic data; (b) simulation & algorithm training enabled by distributed multi-agent architecture and human-AI collaboration mechanism; (c) hardware integration & testing utilizing ROS compatibility for direct validation of autonomous driving algorithms on physical platforms.
III Sky-Drive Workflow
III-A Overview
The workflow begins with the DT framework, which creates high-fidelity virtual replicas of transportation systems through multi-source data integration. These virtual environments feed into the distributed multi-agent architecture, enabling synchronized simulations across multiple devices and supporting complex interactions between autonomous agents. The simulation environment created by these two modules serves as the testing ground for the multi-modal human-in-the-loop framework, which captures comprehensive behavioral data from human participants. This data is then processed and utilized by the human-AI collaboration mechanism, facilitating knowledge exchange between humans and autonomous systems. The foundation models integration will enhance system and agent-level capabilities, enabling global observations for performance feedback and aiding individual agents in better understanding human behavior patterns. Finally, the HIL framework connects with the DT framework, enabling real-world algorithm validation while feeding real-world performance data back into the simulation.
III-B Details
Sky-Drive’s detailed workflow, shown in Fig. 2, consists of three primary stages that form a continuous feedback loop:
III-B1 Scenario Generation & Data Collection
As shown in Fig. 2 (a), this stage employs two complementary approaches to ensure comprehensive scenario coverage: (i) Sky-Drive leverages CARLA and Unreal Engine to generate customizable urban environments with detailed road networks, traffic rules, and environmental conditions, enabling controlled testing of specific driving scenarios. (ii) The DT framework imports real-world data through multi-source integration, including high-precision maps collected by lab-developed AVs, open-source data, and real-world traffic data collection. This data undergoes sophisticated categorization and twinning processes to create digital replicas of physical environments.
III-B2 Simulation & Algorithm Training
As shown in Fig. 2 (b), this stage processes the generated scenarios through an integrated learning pipeline with four interconnected components: (i) The distributed multi-agent architecture enables the concurrent operation of multiple agents across different terminals, facilitating complex traffic interactions in a shared environment while maintaining synchronization. (ii) The human-in-the-loop component integrates multiple human participants, capturing human behavior through immersive interfaces and allowing for real-time feedback collection. (iii) Sky-Drive will integrate LLMs to enhance simulation capabilities, facilitating natural communication between human participants and autonomous systems for more intuitive interaction and knowledge transfer. (iv) The human-AI collaboration mechanism integrates human feedback and domain knowledge into AI training, creating a continuous learning loop where humans inform AI systems and AI provides feedback to humans.
III-B3 Hardware Integration & Testing
As illustrated in Fig. 2 (c), this stage bridges simulation and physical deployment through two key components: (i) While the full HIL framework is planned for future development, the current architecture already supports connections to external hardware through standardized ROS interfaces. The lab-developed Ford E-transit electric van serves as the primary testbed, equipped with dashboard monitors, power modules, and a computing rack for algorithm deployment. Testing is primarily conducted at the Madison College Public Safety Training Center in Columbus, WI, and MGA Research Corporation in Burlington, WI. (ii) Sky-Drive also supports testing of vehicle-to-everything (V2X) communication protocols using Cohda Wireless devices, enabling the evaluation of cooperative perception and decision-making across multiple vehicles and infrastructure elements. This testing is crucial for validating the performance of autonomous systems in complex traffic environments.
III-C Case Demonstration
To demonstrate Sky-Drive’s workflow, consider the case of personalized autonomous driving. In this case, Sky-Drive develops an autonomous system that tailors its behavior to the driver’s unique preferences, learning from their driving style and comfort levels.
The workflow begins with the DT framework, which creates high-fidelity virtual replicas of real-world traffic environments using data such as high-precision maps. These environments are then input into the distributed multi-agent architecture, enabling simulations of complex interactions between AVs and humans. During the simulation and training stage, real-time feedback from the driver, such as “It’s too fast” or “The acceleration is too harsh,” is processed by LLMs to infer preferences regarding acceleration and driving style. This feedback is integrated into the human-AI collaboration mechanism, forming a continuous learning loop where the system adapts its driving strategies and provides more personalized guidance. The HIL framework connects the system to physical platforms, validating the personalized driving algorithm in real-world scenarios. This closed-loop workflow enables the development and validation of personalized autonomous systems, from concept testing to real-world deployment, ensuring safety and reliability.
IV Sky-Drive Features
Figure 3: Illustration of Sky-Drive’s distributed multi-agent architecture. Sky-Drive enables synchronized simulation across multiple terminals while maintaining precise real-time interactions between AVs, HVs, and pedestrians through a sophisticated RPC networking model and Socket.IO-based communication platform, supporting comprehensive data collection and real-time analysis of multi-agent behaviors.
IV-A Distributed Multi-agent Architecture
Future transportation systems will consist of multiple intelligent agents, such as AVs, HVs, and pedestrians, each operating independently, requiring simulation systems that can model and control these agents separately while ensuring seamless interaction between them. Existing platforms, such as Nocturne [20], MetaDrive [14], and Waymax [18], primarily focus on simplified multi-agent interactions on a single machine, limiting their ability to model such complexity. To address this, Sky-Drive introduces a novel distributed multi-agent architecture that enables the synchronized simulation of multiple independently operating agents across different computing devices.
IV-A1 System Architecture
At the core of Sky-Drive lies a sophisticated RPC networking model built upon CARLA, using the rpclib library. This extension enhances CARLA’s vehicle control system, enabling crucial improvements for distributed multi-agent simulation. As shown in Fig. 3 (c), Terminal 1 functions as the host (server) that maintains the global simulation environment, while Terminals 2-4 act as clients controlling different agent types. Each terminal independently controls its corresponding agent through various input devices, while ensuring seamless interaction with other agents in the shared environment. The host terminal is responsible for scene customization and map generation, which is then distributed to the client terminals. The scene generation component (Fig. 3 (a)) creates detailed virtual environments with customizable traffic conditions, weather patterns, and road infrastructure, supporting multiple agent types, including AI-controlled AVs, HVs, pedestrians controlled via VR, and rule-based AVs following predefined behaviors (Fig. 1 (a)).
IV-A2 Communication Infrastructure
The communication infrastructure employs a dual-port TCP system on each terminal, enabling robust bidirectional data exchange between the host and clients. Sky-Drive’s hybrid networking approach ensures optimal performance. For time-critical operations, Sky-Drive utilize a dedicated local area network (LAN) with high-performance switches and Ethernet connections, achieving low latency of 0.3 milliseconds for smooth real-time interactions among agents. For scenarios requiring broader network coverage, such as geographically distributed research across Purdue University and University of Wisconsin-Madison (Fig. 3 (d)), virtual LAN (VLAN) configurations extend the platform’s reach while maintaining communication efficiency.
IV-A3 Real-time Monitoring Platform
A key component of Sky-Drive’s distributed architecture is its real-time monitoring and data management system. Complementing the core networking infrastructure, Sky-Drive has developed a Socket.IO-based communication platform that tracks agent data, including position coordinates, velocity metrics, live video feeds, and sensor readings. As shown in Fig. 3 (b), the platform features a web-based system that provides real-time visualization of agent activities. It streams data to a centralized system where agent interactions are monitored and analyzed in real-time. All simulation data, including agent states, environmental conditions, and interaction events, are logged in a centralized database, enabling comprehensive post-simulation analysis and scenario reproduction.
IV-B Multi-modal Human-in-the-loop Framework
To capture human preferences and cognitive states for adaptive AI behavior, Sky-Drive develops a multi-modal human-in-the-loop framework, illustrated in Fig. 1 (b), that collects and synchronizes gaze patterns, voice commands, facial expressions, physiological signals, and control actions across multiple modalities.
IV-B1 Eye Tracking
Sky-Drive provides an immersive experience through a custom-developed VR interface built on top of the Unreal Engine. Participants engage in the simulation using an HTC Vive Pro Eye headset, which supports full six degrees of freedom (6-DoF) head tracking via SteamVR and integrated eye tracking via the SRanipal SDK. The system captures high-frequency (up to 120 Hz) behavioral signals, including 3D gaze vectors, pupil positions and diameters, eye openness, and fixation points. These signals are critical for analyzing driver attention distribution, situational awareness, and cognitive workload during complex driving tasks.
IV-B2 Voice Interaction
Sky-Drive supports voice commands as an explicit behavioral input modality. Spoken language is transcribed via Whisper, an OpenAI automatic speech recognition (ASR) model [55], and then interpreted by LLMs. The system extracts driver intent and sentiment from structured commands (“slow down at the next intersection”) and informal feedback (“too fast”), mapping them into semantic driving directives or policy preferences to guide AI behavior.
IV-B3 Facial Expression Recognition
A high-resolution in-cabin camera captures facial micro-expressions in real time. Sky-Drive employs expression classification models trained on affective datasets to recognize expressions such as stress, confusion, or satisfaction. These cues serve as implicit indicators of driver state and comfort, enabling real-time adaptation of AI behavior and intervention when necessary.
IV-B4 Physiological Signal Monitoring
Physiological states such as stress and alertness are inferred through biometric signals collected by wearable devices. Sky-Drive integrates the Garmin vívoactive 5 smartwatch to continuously monitor heart rate and heart rate variability (HRV). These physiological signals are synchronized with other behavioral data streams, providing additional channels to model driver arousal, cognitive workload, and fatigue.
IV-B5 Steer Wheel
The ego vehicle is equipped with a Logitech G920 racing wheel and pedal system, with force feedback enabled through the open-source Logitech Wheel Plugin. Steering, throttle, braking, and signaling inputs are logged in parallel with gaze and head pose data. This setup supports realistic driving control and is fully compatible with CARLA’s ScenarioRunner for scenario-based experiments.
IV-C Human-AI Collaboration Mechanism
Sky-Drive implements an adaptive human-AI collaboration mechanism that enables continuous, bidirectional knowledge exchange between human users and AI-enabled autonomous systems. As shown in Fig.1 (d), this mechanism is built on two complementary modules: HAIM and AIHM.
IV-C1 Human as AI Mentor
In the HAIM, humans serve as real-time mentors to AVs, guiding AI learning through two key sources of human knowledge: (i) Individual behavioral knowledge, encompassing both explicit behaviors (e.g., takeovers, voice commands, touchscreen interactions) and implicit signals (e.g., facial expressions, eye movements, physiological responses), captured via Sky-Drive’s multi-modal interface [8, 56]; (ii) Domain knowledge from transportation science, including established models such as IDM and MOBIL that encode long-standing rules of human driving behavior [57].
The HAIM adopts an RL paradigm enhanced by human preference modeling and physics-informed priors to incorporate this dual-source knowledge. Rather than relying on handcrafted reward functions, the HAIM formulates learning as preference-based policy optimization. Frequent human takeovers in specific contexts (e.g., intersections, merges) are treated as implicit indicators of policy failure, shaping cost signals or trajectory ranking. Meanwhile, physics-based models act as behavioral constraints to ensure learned policies remain safe, interpretable, and socially compliant. This design improves sample efficiency, reduces unsafe exploration, and fosters human trust in the AI system.
IV-C2 AI as Human Mentor
In parallel, the AIHM enables AI to function as a real-time coach for human drivers. It leverages Physics-Enhanced Residual Learning (PERL) [58, 59] to generate optimal driving paths that consider vehicle dynamics, safety margins, and individual driving styles [60]. These reference trajectories are visualized in real time via VR or in-vehicle displays and are continuously updated based on driver performance. AIHM evaluates drivers using metrics such as path deviation, response latency, control stability, and situational awareness. Personalized feedback is delivered through annotated replays, visual heatmaps, and AI-generated verbal summaries.
A key innovation of AIHM is the use of generative AI for scenario customization [61]. Based on performance analytics, the system dynamically generates targeted training tasks—such as emergency stops or lane changes—to address specific weaknesses. The level of guidance is continuously adjusted using real-time physiological and behavioral signals: when elevated stress levels (e.g., increased heart rate, frequent steering corrections) are detected, the system reduces scenario complexity and provides calming feedback. Conversely, as the driver demonstrates improved performance, the system introduces more challenging conditions to encourage continued skill development [62].
IV-D Digital Twin Framework
AI algorithms trained in simulation often fail to generalize to real-world traffic due to the lack of environmental fidelity. To address this sim-to-real gap and ensure practical applicability, Sky-Drive introduces a DT framework that creates dynamic, high-fidelity replicas of real transportation systems.
As illustrated in Fig.2 (a), the DT framework consists of two core components: data integration and virtual environment construction. The multi-source data integration layer fuses static and real-time inputs from traffic cameras, loop detectors, connected vehicle telemetry, GPS traces, historical traffic records, and high-definition maps collected using lab-developed AVs equipped with LiDAR and radar. These inputs undergo temporal alignment, spatial correlation, and feature extraction to ensure semantic consistency across sources.
The virtual environment is built on CARLA and Unreal Engine and integrates real-time sensor data and computer vision models to detect and track road users for both rendering and trajectory prediction. By employing video recognition and object tracking models, the system reconstructs road user trajectories and maps them into the digital replica, enabling visual analytics, risk prediction, and event replay. Sky-Drive has implemented a pilot deployment of this framework along the Flex Lane on the Beltline in Dane County, Wisconsin. The DT ingests real-time feeds from WisDOT 511 and historical records from WisTransPortal, enabling dynamic reconstruction of traffic states and generation of predictive insights.
V Sky-Drive Application Case
Figure 4: VR-based experimental setup for studying VRU-AV interactions at unsignalized intersections4 .
V-A VR-based VRU-AV Interaction
Studying interactions between vulnerable road users (VRUs) and AVs is critical for the safe deployment of AV technology in complex urban environments. Although VRU-AV conflicts can lead to serious outcomes, real-world crash data involving these cases remain scarce. More importantly, collecting such data in real traffic is unsafe, difficult to reproduce, and often restricted by ethical constraints. To address this challenge, Sky-Drive provides a VR-enabled platform for investigating VRU-AV interactions in a controlled, immersive, and data-rich environment. Its distributed multi-agent simulation architecture enables synchronized control of multiple agents—across separate terminals and devices—while maintaining real-time coordination. This setup is particularly valuable for modeling high-risk scenarios that are difficult to observe or replicate in the physical world.
As shown in Fig. 4, we conducted a case study focused on right-turn conflicts at unsignalized intersections—a scenario frequently associated with accidents in urban environments. This study leveraged Sky-Drive’s synchronized multi-terminal architecture in a novel experimental setup where human participants experienced the scenario from the pedestrian’s perspective through immersive VR, while researchers controlled an AV making right turns from a separate terminal. During each interaction, Sky-Drive captured multimodal behavioral data from both the AV and the pedestrian. The VR recorded 3D gaze vectors, eye fixations, and reaction times from the pedestrian, while simultaneously logging control signals, deceleration profiles, and trajectory predictions from the AV.
This configuration allows researchers to analyze both the physical outcomes (e.g., successful yielding, near-misses, pedestrian hesitation) and the cognitive-emotional states of the human participant, offering insight into how VRUs perceive and respond to AV behavior.
V-B HAIM-based Deep Reinforcement Learning
To validate the HAIM module, we implemented and tested HAIM-DRL [8], a reward-free RL approach that enables AI agents to learn driving behavior directly from human interventions. This demonstration serves as a proof-of-concept for the HAIM module’s core functionality—leveraging real-time human feedback to guide policy learning—within the multi-agent, simulation-rich environment of Sky-Drive.
Sky-Drive enables HAIM-DRL by detecting and recording steering takeovers, synchronized with vehicle state and surrounding scene context. Within its multi-agent traffic simulation environment, human participants intervene when dissatisfied with the AV’s behavior (e.g., aggressive merging, unsafe following), implicitly indicating suboptimal actions. These interventions are used to construct preference comparisons between pre- and post-takeover trajectories, allowing the agent to identify and avoid human-disapproved actions and iteratively refine its driving policy [8].
Mathematically, the HAIM-DRL can be defined as follows:
$\pi^{*}_{\text{AV}}=\arg\min_{\pi_{\text{AV}}}\mathbb{E}_{s_{t}\sim d_{\pi_{\text{AV}}}}\left[\mathcal{L}\left(\pi_{\text{AV}}(\cdot\mid s_{t}),\pi_{\text{human}}(\cdot\mid s_{t})\right)\right],$ (1)
where $d_{\pi_{\text{AV}}}$ represents the state distribution induced by the agent’s policy $\pi_{\text{AV}}$, and $\mathcal{L}(\cdot,\cdot)$ is a measure of discrepancy (e.g., KL divergence). By minimizing this discrepancy over the state distribution, the AI agent is encouraged to align its behavior with human preferences.
The actual trajectory during the training process is determined by the mixed behavior policy:
$\pi_{\text{mix}}(a\mid s)=\pi_{\text{AV}}(a\mid s)(1-I(s,a))+\pi_{\text{human}}(a\mid s)F(s)$ (2)
[…]
The overall learning objective of HAIM-DRL is specifically designed as [8]:
$\max_{\pi}\;\mathbb{E}\big{[}$ $\psi\hat{Q}(s_{t},a_{t}^{\text{AV}})-\alpha\log\pi_{\text{AV}}(a_{t}^{\text{AV}}\mid s_{t};\theta)$ (3)
$-\beta Q^{\text{EX}}(s_{t},a_{t}^{\text{AV}})-\varphi Q^{\text{IM}}(s_{t},a_{t}^{\text{AV}})\;\big{]}.$
In the Eq. (3), the first term guides the agent to align with human-preferred behavior by minimizing the value discrepancy between its own actions and those demonstrated by the human mentor. The second term introduces an entropy regularization factor that encourages the agent to explore diverse strategies. The third term penalizes actions that frequently trigger human takeovers. The fourth term constrains the agent to minimize disturbances to surrounding traffic.
TABLE II: The performance of PPO, HACO, and HAIM-DRL methods.
Method Test Safety Violation Test Return Test Disturbance Rate Test Success Rate Train Samples
PPO 80.84 1591.00 - 0.35 500,000
HACO 12.14 1578.43 0.0137 0.35 8,000
HAIM-DRL 11.25 1590.85 0.0121 0.38 8,000
Note: The results are based on data reported in [8]. For detailed definitions of evaluation metrics and descriptions of baseline methods, please refer to the original paper.
As evidenced by Tab. II, the HAIM-DRL was successfully implemented and evaluated within the Sky-Drive platform, demonstrating clear advantages over conventional RL methods. Compared with PPO, HAIM-DRL achieves a drastic reduction in safety violations and eliminates the need for large-scale training data, reaching comparable or superior performance with only 8,000 samples. Compared with HACO, which also leverages human interventions, HAIM-DRL further improves test return, reduces disturbance rate, and increases the success rate from 0.35 to 0.38. These results validate Sky-Drive’s capability to support closed-loop human-AI training, enabling efficient, human-aligned policy learning through real-time feedback and preference-driven optimization.
TABLE III: Performance comparison of PE-RLHF with different physics-based model combinations.
Method Driving Operation Training Testing
Total Safety Violation $\downarrow$ Episodic Return $\uparrow$ Success Rate (%) $\uparrow$ Safety Violation $\downarrow$ Travel Distance $\uparrow$ Travel Velocity $\uparrow$ Total Overtake Count $\uparrow$
IDM-MOBIL Longitudinal & Lateral - 206.30 $\pm$35.23 0.31 $\pm$0.15 0.49 $\pm$0.08 108.56 $\pm$55.23 19.78 $\pm$2.67 0 $\pm$0
PE-RLHF (without) - 39.45 $\pm$ 12.32 302.67 $\pm$ 21.88 0.73 $\pm$ 0.05 1.48 $\pm$ 0.43 138.23 $\pm$ 4.28 16.58 $\pm$ 0.96 6.14 $\pm$ 1.12
PE-RLHF (with IDM) Longitudinal only 28.79 $\pm$ 9.97 348.52 $\pm$ 19.67 0.79 $\pm$ 0.03 0.98 $\pm$ 0.29 149.87 $\pm$ 4.10 18.92 $\pm$ 0.94 7.83 $\pm$ 1.03
PE-RLHF (with MOBIL) Lateral only 21.56 $\pm$ 8.54 368.11 $\pm$ 18.45 0.81 $\pm$ 0.04 0.74 $\pm$ 0.19 159.34 $\pm$ 3.14 20.43 $\pm$ 0.51 9.76 $\pm$ 1.17
PE-RLHF (with IDM-MOBIL) Longitudinal & Lateral 16.61 $\pm$ 9.96 391.48 $\pm$ 20.47 0.85 $\pm$ 0.04 0.47 $\pm$ 0.01 177.00 $\pm$ 3.74 21.85 $\pm$ 0.02 16.33 $\pm$ 4.61
Note: The best results are marked in bold. The results are based on data reported in [57]. For detailed definitions of evaluation metrics and descriptions of baseline methods, please refer to the original paper.
V-C Physics-enhanced Reinforcement Learning with Human Feedback
To validate the integration of knowledge in transportation science within Sky-Drive, we implemented and tested the physics-enhanced reinforcement learning with human feedback (PE-RLHF) [57]. Unlike traditional methods that may falter with imperfect human feedback, PE-RLHF establishes a trustworthy safety performance lower bound through well-established traffic flow models.
PE-RLHF implements this idea through three policies: a human policy $\pi_{\text{human}}$, a physics-based policy $\pi_{\text{phy}}$ derived from traffic flow models, and a learning agent policy $\pi_{\text{AV}}$. When no takeover occurs, the AV agent executes $\pi_{\text{AV}}$ and updates its policy through environment-driven exploration. When human intervention is detected, the PE-RLHF compares the expected values of actions suggested by the human and the physics-based policy.
$\mathcal{T}_{\text{PE-HAI}}(s)=\begin{cases}a_{\text{hybrid}},&\text{if takeover}\\ a_{\text{AV}},&\text{otherwise}\end{cases}$ (4)
In detail, if the human action is expected to yield better outcomes, it is selected; otherwise, the physics-based action is executed. This hybrid policy guarantees that decisions adhere to a minimum standard of safety and efficiency:
$a_{\text{hybrid}}$ $=T_{\text{select}}(s)$ (5)
=$=\begin{cases}a_{\text{human}},&\text{if }Q^{\phi}(s,a)-Q^{\phi}(s,a)\geq\varepsilon_{\text{select}}\\ a_{\text{phy}},&\text{otherwise}\end{cases}$
The Eq. (5) ensures that the system always executes the action with higher expected value, establishing a performance floor guaranteed by interpretable physics-based models, even when human feedback quality deteriorates.
The overall learning objective of PE-RLHF is formulated as [57]:
$\max_{\pi}\mathbb{E}\left[\psi\hat{Q}(s_{t},a_{t}^{\text{hybrid}})-\alpha\log\pi_{\text{AV}}(a_{t}^{\text{AV}}|s_{t};\theta)-\beta Q^{\text{int}}(s_{t},a_{t}^{\text{AV}})\right]$ (6)
where $\hat{Q}(s_{t},a_{t}^{\text{hybrid}})$ is a proxy value function, the entropy term encourages exploration, and $Q^{\text{int}}(s_{t},a_{t}^{\text{AV}})$ minimizes the need for human intervention.
Tab. III presents the performance comparison of PE-RLHF under different configurations, including standalone physics-based control (IDM-MOBIL) and PE-RLHF variants with or without integrated physics models. The full PE-RLHF configuration consistently outperforms all baselines across all stages. PE-RLHF improves episodic return and success rate by nearly 90% compared to the standalone model. Meanwhile, it achieves the lowest safety violation (0.47) and the longest travel distance (177.00 m). Moreover, the full configuration reaches the highest travel velocity (21.85 km/h) and completes the most overtaking maneuvers (16.33), whereas the IDM-MOBIL baseline fails to overtake at all. These results validate Sky-Drive’s ability to support the integration of physics knowledge and human feedback for learning safe and efficient autonomous driving policies.
V-D Vision Language Model-Enabled Reinforcement Learning
TABLE IV: Performance comparison with baselines during testing. Mean and standard deviation over 3 seeds.
Model Average Speed $\uparrow$ Route Completion $\uparrow$ Traveled Distance $\uparrow$ Collision Rate $\downarrow$ Success Rate $\uparrow$
VLM-SR 0.53 $\pm$ 0.27 0.02 $\pm$ 0.00 47.9 $\pm$ 9.2 0.18 $\pm$ 0.25 0.0 $\pm$ 0.0
RoboCLIP 0.44 $\pm$ 0.05 0.07 $\pm$ 0.03 146.3 $\pm$ 62.3 1.05 $\pm$ 0.58 0.0 $\pm$ 0.0
VLM-RM 0.20 $\pm$ 0.05 0.02 $\pm$ 0.01 35.9 $\pm$ 25.8 0.003 $\pm$ 0.005 0.0 $\pm$ 0.0
LORD 0.17 $\pm$ 0.08 0.02 $\pm$ 0.02 45.1 $\pm$ 57.1 0.02 $\pm$ 0.02 0.0 $\pm$ 0.0
LORD-Speed 18.9 $\pm$ 0.36 0.87 $\pm$ 0.08 1783.4 $\pm$ 172.8 2.80 $\pm$ 1.16 0.67 $\pm$ 0.05
VLM-RL (ours) 19.3 $\pm$ 1.29 0.97 $\pm$ 0.03 2028.2 $\pm$ 96.6 0.02 $\pm$ 0.03 0.93 $\pm$ 0.04
Note: The best results are marked in bold. The results are based on data reported in [56]. For detailed definitions of evaluation metrics and descriptions of baseline methods, please refer to the original paper.
To validate Sky-Drive’s capability to support VLM-enabled RL, we implemented the VLM-RL [56], which integrates pre-trained VLMs with RL to generate semantic reward signals from image observations and natural language goals. This demonstration showcases Sky-Drive’s ability to enable high-level, human-interpretable guidance for safe and efficient autonomous driving.
At the core of VLM-RL is the contrasting language goal (CLG)-as-reward paradigm, which uses pre-trained VLMs to compute semantic similarity between driving states and paired language descriptions. Positive goals (e.g., “the road is clear with no accidents”) and negative goals (e.g., “two cars have collided”) are used to guide the agent’s behavior by comparing how closely its current state aligns with each description. The reward is computed by encoding visual input via CLIP’s image encoder and goals via its text encoder, both mapped into a shared latent space [56]:
$R_{\text{CLG}}(s)=$ $\alpha\cdot\text{sim}(\text{VLM}_{\text{I}}(\psi(s)),\text{VLM}_{\text{L}}(l_{\text{pos}}))$ (7)
$-\beta\cdot\text{sim}(\text{VLM}_{\text{I}}(\psi(s)),\text{VLM}_{\text{L}}(l_{\text{neg}}))$
where $l_{\text{pos}}$ and $l_{\text{neg}}$ are the positive and negative language goals, $\text{VLM}_{\text{I}}$ and $\text{VLM}_{\text{L}}$ denote the image and language encoders of the pre-trained VLM, $\psi(s)$ is the visual preprocessing function, and $\text{sim}(\cdot,\cdot)$ represents cosine similarity. The weights $\alpha$ and $\beta$ control the influence of the positive and negative goals, respectively.
To improve reward stability, VLM-RL introduces a hierarchical reward synthesis strategy that combines CLG-based semantic rewards with low-level vehicle state signals such as speed alignment, lane deviation, and directional consistency. The synthesized reward is defined as [56]:
$R_{\text{synthesis}}(s)=r_{\text{speed}}(s)\cdot f_{\text{center}}(s)\cdot f_{\text{angle}}(s)\cdot f_{\text{stability}}(s)$ (8)
where $r_{\text{speed}}(s)=1-\frac{|v-v_{\text{target}}|}{v_{\max}}$ measures speed alignment with respect to the target velocity $v_{\text{target}}=r_{t}^{\prime\text{CLG}}\cdot v_{\max}$; $f_{\text{center}}(s)$ evaluates the vehicle’s lateral position relative to the lane center; $f_{\text{angle}}(s)$ reflects the vehicle’s orientation with respect to the road direction; and $f_{\text{stability}}(s)$ quantifies the temporal consistency of the vehicle’s lateral positioning.
As shown in Tab. IV, VLM-RL significantly outperforms existing approaches across all key metrics. VLM-RL achieves the highest success rate and route completion, while maintaining a low collision speed of 0.02 km/h—matching the safety level of the most conservative baselines. Unlike existing VLM-based methods, which suffer from overly cautious behavior and near-zero task success, VLM-RL balances safety with efficiency, reaching an average speed of 19.3 km/h and a total driving distance of 2028.2 m. Compared to strong LLM-based methods such as Revolve, VLM-RL maintains comparable success and completion rates while drastically reducing collision speed. The successful implementation of VLM-RL within the Sky-Drive platform validates its capability to support large-scale, multimodal policy learning grounded in human-understandable semantics.
V-E Personalized Safety-Critical Curriculum Learning
TABLE V: Performance comparison with baselines in the safety-critical test scenarios.
Model Episode Reward $\uparrow$ Road Completion (%) $\uparrow$ Total Distance $\uparrow$ Crash Rate (%) $\downarrow$ Average Speed $\uparrow$ Failure-to- Success Rate (%) $\uparrow$ Success-to- Success Rate (%) $\uparrow$
SAC 38.4 $\pm$ 1.97 63.2 $\pm$ 1.21 40.9 $\pm$ 1.34 30.5 $\pm$ 2.33 9.25 $\pm$ 0.07 30.4 $\pm$ 7.00 56.9 $\pm$ 15.1
PPO 38.4 $\pm$ 0.86 62.7 $\pm$ 1.05 40.0 $\pm$ 0.70 32.0 $\pm$ 2.02 9.94 $\pm$ 0.30 26.7 $\pm$ 0.89 41.7 $\pm$ 8.33
TD3 42.4 $\pm$ 1.01 65.2 $\pm$ 1.40 42.6 $\pm$ 1.26 39.7 $\pm$ 1.04 8.02 $\pm$ 0.77 28.6 $\pm$ 2.79 64.3 $\pm$ 21.4
CAT 42.5 $\pm$ 3.95 66.6 $\pm$ 4.37 43.4 $\pm$ 3.48 32.1 $\pm$ 2.08 8.36 $\pm$ 1.17 35.2 $\pm$ 3.44 67.5 $\pm$ 7.50
CLIC 39.3 $\pm$ 0.72 64.3 $\pm$ 0.40 41.6 $\pm$ 0.78 26.2 $\pm$ 1.17 9.21 $\pm$ 0.26 34.7 $\pm$ 2.67 61.9 $\pm$ 26.9
CurricuVLM (ours) 48.9 $\pm$ 1.53 73.4 $\pm$ 1.66 48.4 $\pm$ 1.31 25.1 $\pm$ 1.17 9.45 $\pm$ 0.16 39.1 $\pm$ 0.66 73.5 $\pm$ 21.1
Note: The best results are marked in bold. The results are based on data reported in [62]. For detailed definitions of evaluation metrics and descriptions of baseline methods, please refer to the original paper.
To validate Sky-Drive’s capability to support adaptive scenario generation and curriculum learning, we implement the CurricuVLM [62]. CurricuVLM integrates VLMs to enable personalized, safety-critical training scenarios tailored to the evolving weaknesses of autonomous driving agents.
The core innovation of CurricuVLM lies in bridging the gap between scenario generation and policy learning. By continuously monitoring agent performance, the framework identifies failure patterns through a two-stage behavior analysis pipeline: VLMs are first used to extract rich visual descriptions of unsafe events, which are then interpreted by a GPT-4o-based analyzer to uncover behavioral limitations. This process enables semantic understanding of critical driving mistakes without manual annotation.
Based on the analysis, scenario generation is formulated as a conditional trajectory generation problem:
$P(Y^{AV},Y^{BV}|I,X)$ (9)
where $X$ encodes historical context (e.g., maps, past trajectories), $I$ contains semantic insights from behavior analysis, and $Y^{AV}$, $Y^{BV}$ denote future trajectories of the ego and background vehicles, respectively.
The framework optimizes $Y^{BV}$ to generate targeted, informative interactions via:
$Y^{\text{BV}*}=$ $\arg\max_{Y^{\text{BV}}}\;P(Y^{\text{BV}}\mid X)\cdot$ (10)
$\sum_{Y^{\text{AV}}\sim\mathcal{Y}(\pi)}P(Y^{\text{AV}}\mid Y^{\text{BV}},X)\cdot P(I\mid Y^{\text{AV}},Y^{\text{BV}})$
where $\mathcal{Y}(\pi)$ denotes the trajectory distribution induced by the current policy $\pi$, and $P(I\mid Y^{\text{AV}},Y^{\text{BV}})$ measures how well the generated scenario aligns with the identified behavioral insight. This formulation encourages the background vehicle behavior to induce targeted policy responses from the AV agent, forming the foundation for automated curriculum construction.
As shown in Tab. V, CurricuVLM achieves the best overall performance across all key metrics, demonstrating both high safety and training effectiveness. In terms of task performance, CurricuVLM achieves the highest episode reward (48.9) and road completion rate (73.4%), while maintaining a low crash rate (25.1%), outperforming strong baselines such as CAT and CLIC. It also records the highest total driving distance (48.4m) and failure-to-success rate (39.1%), indicating superior adaptability to previously failed scenarios. Meanwhile, its success-to-success rate (73.5%) reflects strong behavioral consistency and learning stability. These results validate that CurricuVLM not only enhances policy robustness under long-tail safety-critical scenarios, but also integrates seamlessly into Sky-Drive’s human-AI collaboration mechanism.
V-F Interactive and Editable Traffic Scenario Generation
Figure 5: Qualitative examples. Each scenario is downsampled to four frames for visualisation.
To validate Sky-Drive’s capability in supporting multimodal, human-centered scenario generation, we implemented the Talk2Traffic [61], which enables intuitive and editable traffic scenario creation through natural language, speech, and sketch-based inputs, as illustrated in Fig. 5. It demonstrates Sky-Drive’s support for seamless integration of multimodal large language models (MLLMs) into simulation workflows for autonomous driving research.
The core idea of Talk2Traffic is to bridge the gap between human designers’ intent and executable traffic simulation. The system first interprets user instructions through a multimodal encoder [61]:
$\mathbf{z}=\text{MLLM}(p,l,s),$ (11)
where $p$ is the task prompt, $l$ represents textual or spoken language input, and $s$ denotes sketch-based spatial constraints. The output $\mathbf{z}$ is a structured scene representation capturing key elements such as map layout, agent behaviors, and environmental conditions.
To generate executable code, Talk2Traffic adopts a retrieval-augmented generation (RAG) strategy based on a curated database of description-code pairs:
$\mathcal{D}=\{(d_{j},c_{j})|j\in\{1,\ldots,m\}\},$ (12)
where $d_{j}$ is a natural language description and $c_{j}$ is the corresponding Scenic code snippet. Relevant snippets are retrieved using semantic similarity. A key feature of Talk2Traffic is its interactive refinement mechanism, where users can iteratively modify the scenario based on feedback.
Experimental results demonstrate Talk2Traffic’s superior capability in generating diverse and challenging scenarios. It achieves the lowest average collision rate (0.877) across all scenario types, surpassing the next-best baseline by 4.6%. In high-complexity interactions such as red light running (0.900) and unprotected left turn (0.833), Talk2Traffic consistently outperforms other methods. Some qualitative examples of generated scenarios are illustrated in Fig. 5. By integrating Talk2Traffic into the AIHM module, Sky-Drive enables dynamic and goal-directed scenario generation, allowing personalized training and evaluation of autonomous agents under richly varied, realistic, and safety-critical conditions.
V-G Accident Data Replay
Figure 6: Accident data replay framework for systematic traffic incident analysis.
To validate Sky-Drive’s capability in supporting real-world accident reconstruction and analysis, we implemented an accident data replay framework that enables systematic reproduction of traffic collisions within Sky-Drive environment. This framework addresses a fundamental need in autonomous driving development: understanding and learning from real accidents in a safe, repeatable, and controlled setting.
The replay pipeline centers around CenterTrack [63], an advanced multi-object tracking algorithm used to extract object trajectories from raw accident video footage. These 2D trajectories are then mapped into 3D space and replayed in Unreal Engine through Sky-Drive’s integrated simulation backend. As shown in Fig. 6, the reconstructed scenes preserve key dynamics such as vehicle positions, speeds, and interaction sequences, along with contextual factors like road layout and weather. Specially, Sky-Drive incorporates a robust reconstruction validation process to ensure fidelity. A procedural matching algorithm selects the most appropriate simulation maps based on road topology and scene semantics. A built-in quality assessment module scores the visual and kinematic similarity between the replayed and original sequences, flagging low-fidelity cases for refinement. The framework also supports unsupervised domain adaptation to improve trajectory accuracy, while offering manual editing tools when needed to ensure precise alignment with real-world footage.
This replay capability enables several downstream applications: (i) It provides a safe testbed for analyzing accident causation, enabling the development of improved safety mechanisms and behavior prediction models; (ii) It allows RL agents to be trained and evaluated on real-world edge cases, significantly enhancing their robustness in critical scenarios; (iii) It supports regulatory compliance and post-incident investigation by producing detailed, verifiable accident reconstructions. Through this integration, Sky-Drive enables scalable, high-fidelity replay of accident scenarios, positioning itself as a comprehensive platform for evaluating autonomous systems under rare, safety-critical conditions that are otherwise difficult or unsafe to replicate.
VI Future Enhancements
VI-A Foundation Models Integration
VI-A1 Multimodal Behavioral Understanding
Interpreting human behavioral signals in a unified, context-aware manner remains an open challenge. Future iterations of Sky-Drive will leverage LLMs and VLMs to perform cross-modal reasoning across physiological, visual, verbal, and control-based modalities. For example, an elevated heart rate, downward gaze, and a quick verbal cue like “too fast” may collectively indicate the driver’s discomfort with vehicle acceleration. A more nuanced phrase such as “I feel a bit uneasy because the car accelerates too quickly” can be semantically aligned with facial tension and biometric signals like heart rate variability. By combining these signals in the context of traffic density, road geometry, and interaction with nearby vehicles, Sky-Drive can construct rich behavioral profiles far beyond what single-modality systems can achieve. This capability will support personalized feedback generation, trust modeling, and adaptive control within the HAIM and AIHM modules.
VI-A2 Personalized Autonomous Driving
As shown in Fig 7, Sky-Drive will implement a LLM-based system that enables personalized autonomous driving through natural language interaction [64]. Specifically, the system will integrate three core modules: a visual encoder to process real-time camera feeds, an LLM to interpret language inputs, and a route planning module to generate executable commands based on Sky-Drive’s maps. To ensure robustness, Sky-Drive will use a three-stage training pipeline. The first stage uses the BDD-X dataset [65] to align visual and linguistic representations. The second stage fine-tunes language understanding via LoRA techniques on the SDN dataset [66]. The final stage incorporates data generated within the Sky-Drive simulation environment to adapt model responses to realistic driving tasks. This integration allows drivers to provide real-time feedback such as “slow down a bit here” or “take the next left,” and have the vehicle respond accordingly. In the long term, this capability will support personalized, explainable, and user-aligned driving experiences.
Figure 7: LLM-based system enabling personalized autonomous driving.
VI-A3 Traffic Brains
Sky-Drive will position foundation models as intelligent “traffic brains” that govern decision-making in complex, multi-agent traffic environments. While general-purpose models such as Qwen [23], GPT-4 [21], and Llama [22] exhibit strong language and reasoning abilities, they will require domain-specific adaptation to meet the demands of autonomous driving. To address this challenge, Sky-Drive will leverage transportation-specific datasets—including LMDrive [67], CCD [68], DoTA [69], and DriveCoT [70]—to fine-tune pre-trained foundation models. This fine-tuning pipeline is designed to enhance the model’s ability to handle dynamic scenario adaptation, hierarchical reasoning, and multitask decision-making, including generating safe control actions (e.g., steering, throttle, and braking) and predicting critical safety metrics such as time-to-collision (TTC). The refined models will be deployed within the Sky-Drive simulation environment to enable coordinated behavior across AVs and other components, facilitating holistic control and system-level optimization.
VI-B Hardware-in-the-Loop
VI-B1 Simulation-to-Reality Integration
As shown in Fig. 1 (e), the center of the HIL framework is a Ford E-Transit electric van retrofitted with fully automated driving capabilities. The vehicle is equipped with a comprehensive sensor suite—including LiDAR, radar, high-resolution cameras, and OxTS navigation units—and operates on a drive-by-wire system connected to an industrial-grade computing rack. The vehicle’s software stack builds upon ROS-based open-source packages, further augmented with proprietary Sky-Drive modules to support advanced planning, control, and communication tasks. To complement the vehicle-level testing, Sky-Drive also deploys portable roadside infrastructure units outfitted with traffic lights, regulatory signage, cameras, and LiDAR systems. These roadside units enable systematic evaluation of V2X communication and cooperative perception algorithms across diverse environmental settings [71, 72]. By maintaining connectivity with Sky-Drive’s DT environment, the framework allows algorithms to be first validated in simulation and then deployed to physical vehicles with minimal transition cost, significantly accelerating development cycles while ensuring safe real-world performance.
VI-B2 Teleoperated Driving
The HIL framework also establishes a solid foundation for developing and testing teleoperated driving. Teleoperated driving allows humans (teleoperators) to remotely control vehicles, particularly in challenging scenarios, complementing fully/highly autonomous solutions. It is one of the important use cases of V2X communication, specified in the 3GPP standards [73]. Sky-Drive’s ROS integration enables wireless connectivity between its testbed vehicle and human-in-the-loop simulation platform—operated by a teleoperator—via cellular or satellite networks, e.g., 5G. Considering the wild fluctuations of network bandwidth, round-trip time (RTT), jitter time, and packet loss under driving conditions of 5G [74], Sky-Drive facilitates the collaboration between the vehicle and the simulation platform to dynamically decide what data (RGB images, LiDAR point cloud, and/or their pre-processed data) to transmit and how to transmit them to meet the end-to-end latency requirement for the teleoperation, i.e., below 100 milliseconds [75].
VII Conclusions
This paper presented Sky-Drive, a distributed multi-agent simulation platform designed for socially-aware autonomous driving and human-AI collaboration in future transportation systems. Unlike existing simulators that primarily focus on validating single-vehicle performance, Sky-Drive addresses the emerging need to explore complex interactions in mixed traffic environments where various intelligent agents must align with human preferences and societal norms.
Sky-Drive introduces several key innovations: (a) a distributed multi-agent architecture enabling synchronized simulation across multiple terminals, allowing independent agent control while maintaining shared environmental states; (b) a multi-modal human-in-the-loop framework integrating diverse sensors to capture comprehensive behavioral data; (c) a novel human-AI collaboration mechanism to facilitate bidirectional knowledge exchange; and (d) a digital twin framework creating high-fidelity virtual replicas of real transportation systems. The platform’s effectiveness has been demonstrated through multiple application cases, including VR-based vulnerable road user interactions, physics-enhanced reinforcement learning with human feedback, vision-language model-enabled reinforcement learning, personalized curriculum learning, and accident data replay. These applications show Sky-Drive’s potential to advance autonomous driving research beyond traditional metrics of safety and efficiency toward more socially aware and human-aligned behavior.
To further enhance Sky-Drive’s capabilities, we have outlined two major planned functionalities: (i) the integration of foundation models to support multimodal behavior understanding, personalized driving, and system-level optimization via traffic brains; and (ii) a hardware-in-the-loop framework via ROS integration to enable direct validation of algorithms on physical vehicles. These future enhancements will bridge the gap between simulation and reality, ensuring that algorithms are safely evaluated in real-world environments. As autonomous driving technology continues to evolve, Sky-Drive provides a robust platform for ensuring that future transportation systems are not only safe and efficient but also socially aware and aligned with human expectations.